OpenStudy (anonymous):

A coin is tossed 5 times. Find the probability of getting at least 3 heads.

OpenStudy (moneybaby980):

ok

OpenStudy (moneybaby980):

is that a statement or a part of a question

OpenStudy (kropot72):

This can be solved by using the binomial distribution: \[\large P(number\ of\ heads \ge 3)=\left(\begin{matrix}5 \\ 3\end{matrix}\right)0.5^{5}+\left(\begin{matrix}5 \\ 4\end{matrix}\right)0.5^{5}+0.5^{5}\]